A 2.0 kg box is initially at rest on a frictionless surface. A ball is thrown at the side of the box such that the ball is in contact with the box for 0.1 s before bouncing back. Afterwards the box is moving at 1.25 m/s What average force did the ball exert on the box? A) 12 N B) 25 N C) 34 N D) 45 N E) 63 N
Answer #1

We can rewrite Newton’s second law by taking momentum change into consideration as follows:

\sumF = ma

F = mdv / dt

F = dp / dt

-F = m(vf - vi) / dt

-F = m(0 - vi) / dt

F = ( 2 x 1.25) / 0.1

F = 25 N